A large container which has 285 g of water initially at 21 °C is placed in a microwave oven. The
magnetron in the microwave oven can produce 950 W of EM radiation. The microwave is turned
on for 2.5 minutes. Assuming that all the microwave energy is absorbed by the water, how much
water is left in the container?
1
Expert's answer
2019-10-17T11:10:08-0400
"Q=Pt=Mc(T_2-T_1)+(M-m)L"

"(950)(2.5\cdot 60)=(0.285)(4200)(100-21)+(0.285-m)2265000"

"m=0.264\ kg=264\ g"